OpenStudy (anonymous):

Find a polynomial function f(x) with real coefficients that has the given zeros.-1,-1, 2+3i

OpenStudy (anonymous):

Expand \[ (x+1)(x+1) (x- 2 - 3 i)(x-2+3i) \]

OpenStudy (anonymous):

\[ (x+1) (x+1) ((x-2)-3 i) ((x-2)+3 i)=x^4-2 x^3+6 x^2+22 x+13 \]

OpenStudy (anonymous):

If 2 + 3i is a root then 2-3i is also a root

OpenStudy (anonymous):

The polynomial would be in the form; y = (x^2 - 1)(x^2 + bx + c) find b and c. Sum of the 2 roots: (2 + 3i) + (2 - 3i) = 4 , then-> b = -4 Product of 2 roots: c / a = (2 + 3i)*(2 - 3i) = 4 - 9i^2) = (4 + 9) = 13 = Then, c = 13. The trinomial is (x^2 - 4x + 13) Check: D = b^2 - 4ac = -36 -> VD = 6i and -6i 2 roots: x1 = 2 + 3i, and x2 = 2 - 3i. Finally: y = (x^2 - 1)(x^2 - 4x + 13)

OpenStudy (anonymous):

Correction. y = (x + 1)^2*(x^2 - 4x + 13) because the 2 real roots are -1 and -1.